# Homological shifts of a complementary edge ideal

### v1

@@ 4.3. Cycles
We only consider the case when $n=2m+1$ (an odd number), as the proof for the other case is analogous and even simpler. Let $j\in\mathrm{set}(\beta)$. Then $j\in[n-1]$ and $j\in u_{k}$ for some $1\leq k\leq s$. If j=n1=2mj=n-1=2m, then by Lemma 4.8.(3), $\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2m-3},{\bf e}_{2m-1}\}\subseteq\{u_{1},\ldots,u_{s}\}$, which implies $j\in N_{1}(\beta)$. Thus, we may assume $j\leq 2m-1$.
If $j=\min u_{k}$ for some $k\in[s]$, then $j\in\{\min u_{j}:\ 1\leq j\leq s\}\setminus\{n-1\}$. Suppose now that $j\neq\min u_{k}$ for any $k=1,\ldots,s$. By Lemma 4.8, we conclude:
- If $j=2\ell$ for some $\ell$, then $\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-1}\}\subseteq\{u_{1},\ldots,u_{s}\}$;
- If $j=2\ell+1$ for some $\ell$, then $\{{\bf e}_{n},{\bf e}_{2},\ldots,{\bf e}_{2\ell}\}\subseteq\{u_{1},\ldots,u_{s}\}$.
Since $j\leq 2m-1$, we have: If $j=2\ell+1$, then $\ell\leq m-1$; If $j=2\ell$, then $\ell\leq m-1$. This implies $j\in N_{1}(\beta)\cup N_{2}(\beta)$, so the inclusion $\subseteq$ holds. The converse inclusion follows directly from Lemma 4.8.
@@
By the above corollary, it is easy to see that $\mathrm{HS}_{i}(\mathcal{R(}I))$ is generated in degree $\leq n$ for all $i\geq 1$.
Hereafter, we devote to compute the homological shift ideals of powers of the complementary edge ideal of a cycle.
@@
Let $G$ be a cycle of size $n$ and $I$ be its complementary edge ideal. For $1\leq i\leq n-1$ and $s\geq\frac{i}{2}$, the following inclusion holds:
@@
Write $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$ with $u_{j}\in I(G)$ for all $j=1,\ldots,s$. Let $T=\{u_{1},\ldots,u_{s}\}$. Denote
@@
By convention, $\max\emptyset=$-$\infty$. It is clear that $N_{1}(\beta)\cup N_{2}(\beta)=[2\ell_{1}]\cup[2\ell_{2}+1]$. We consider two cases:
@@
(2) $Q\cap\left([2\ell_{1}]\cup[2\ell_{2}+1]\right)=Q_{1}\neq\emptyset$: In this case, we can decompose $Q$ as
@@
where $j_{t}\in f_{t}$ for $t=1,\ldots,r$. These edges $v_{1},\ldots,v_{q},f_{1},\ldots,f_{r}$ are pairwise disjoint, with $2q+r=p$ and $q+r\leq\ell_{2}$. For each $t=1,\ldots,r$, let $i_{t}\in[2\ell_{2}+1]$ be such that $f_{t}=x_{j_{t}}x_{i_{t}}$.
Since $\{v_{1},\ldots,v_{q},f_{1},\ldots,f_{r},{\bf e}_{j_{p+1}},\ldots,{\bf e}_{j_{i}}\}\subseteq\{u_{1},\ldots,u_{s}\}$, we may assume (after reindexing) that
@@
Let $F=\{i_{1},\ldots,i_{r},j_{p+1}+1,\ldots,j_{i}+1\}$, which is an (iq)(i-q)-subset of $[n]$. Thus, ${\bf x}_{Q}\beta\in\frac{\alpha^{i-q}}{{\bf x}_{F}}I^{s-i+q}$, meaning ${\bf x}_{Q}\beta$ belongs to the $k=q$ component of the right-hand side of (4).
The proof for $\ell_{1}>\ell_{2}$ is analogous, so the argument is complete.    ∎
Next, we show that if $i$ is large enough, i.e., $i>\frac{n}{2}$, certain components of the right-hand of (4) may be removed.
@@
If $i\leq\left\lceil\frac{n}{2}\right\rceil-1$, then the right-hand side of (5) coincides with the left-hand side of (4). Thus, the result follows from Lemma 4.11. Suppose now that $i>\left\lceil\frac{n}{2}\right\rceil-1$. We only consider the case where $n=2m$, as the argument for $n=2m+1$ is analogous. We may write i=m1+qi=m-1+q. Note that $q=\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}$. Let $\beta\in\mathcal{G}(I^{s})$ with $i\leq|\mathrm{set}(\beta)|$, and take a subset $Q\subseteq\mathrm{set}(\beta)$ with $|Q|=i$. It suffices to show that ${\bf x}_{Q}\beta$ lies in the right-hand side of (5). Write $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$ with $u_{j}\in I(G)$ for all $j=1,\ldots,s$. Let $T=\{u_{1},\ldots,u_{s}\}$.
@@
$\mathrm{set}(\beta)=[2\ell_{1}]\cup\left(\{\min u_{j}:\ j=1,\ldots,s\}\setminus\{n-1\}\right).$
Note that the edges ${\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2m-3}$ are pairwise disjoint, and their union is the set [2m2][2m-2]. Since $Q\subseteq\mathrm{set}(\beta)\subseteq[2m]$ and |Q|=m1+q|Q|=m-1+q, by the pigeonhole principle, there are $k\geq q$ edges among ${\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2m-3}$ that are entirely contained in $Q$. Let these $k$ edges be ${\bf e}_{2j_{1}-1},{\bf e}_{2j_{2}-1},\ldots,{\bf e}_{2j_{k}-1}$ with $j_{1}<j_{2}<\cdots<j_{k}$.
@@
$\mathrm{set}(\beta)=[2\ell_{2}+1]\cup\left(\{\min u_{j}:\ j=1,\ldots,s\}\setminus\{n-1\}\right).$
@@
Hence, ${\bf x}_{Q}\beta$ belongs to $k=q$ component of the right-hand side of (5). This completes the proof.    ∎
Let us consider the converse inclusion.
@@
Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal. Assume $n\in\{2m,2m+1\}$. Given a subset $F\subseteq[n]$ with $p=|F|$ and an integer $k\geq 0$ such that $k+p<\frac{n}{2}$ (i.e., $k+p\leq m-1$ if $n=2m$, and $k+p\leq m$ if $n=2m+1$), set $i=2k+p$. Then there exist pairwise distinct edges $u_{1},\ldots,u_{k},u_{k+1},\ldots,u_{k+p}$ and a subset $E\subseteq\mathrm{set}\left(\frac{\alpha^{k+p}}{u_{1}\cdots u_{k+p}}\right)$ with $|E|=i$ such that $\frac{\alpha^{k+p}}{{\bf x}_{F}}={\bf x}_{E}\cdot\frac{\alpha^{k+p}}{u_{1}\cdots u_{k+p}}\in\mathrm{HS}_{i}(I^{k+p})$.
@@
We only consider the case $n=2m$; the proof for $n=2m+1$ is analogous. Let $F=\{j_{1}<j_{2}<\cdots<j_{p}\}$. We analyze two cases:
(1) $n\notin F$ (i.e., $j_{p}<n$): Define $t=\max\{a:j_{1}\leq 2k+2,\ j_{2}\leq 2k+4,\ \ldots,\ j_{a}\leq 2k+2a\}$. By convention, $t=0$ if $j_{1}>2k+2$. Then $1\leq j_{1}<\cdots<j_{t}\leq 2k+2t\leq 2m-2$ and $j_{p}>\cdots>j_{t+1}>2k+2t+2$. Set
$\beta=\frac{\alpha^{k+p}}{({\bf e}_{1}{\bf e}_{3}\cdots{\bf e}_{2k+2t-1})({\bf e}_{j_{t+1}-1}\cdots{\bf e}_{j_{p}-1})}.$
Then $E=([2k+2t]\setminus\{j_{1},\ldots,j_{t}\})\cup\{j_{t+1}-1,\ldots,j_{p}-1\}$ is a subset of $\mathrm{set}(\beta)$. Moreover, $|E|=2k+p=i$ and ${\bf x}_{E}\cdot\beta=\frac{\alpha^{k+p}}{{\bf x}_{F}}$, as required.
(2) $n\in F$ (i.e., $j_{p}=n$): As in case (1), there exists $t$ with $0\leq t\leq p-1$ such that $\{j_{1},\ldots,j_{t}\}\subseteq[2k+2t+1]\subseteq[2m-1]$ but $j_{t+1}>2k+2t+3$. Define
$\beta=\frac{\alpha^{k+p}}{({\bf e}_{0}{\bf e}_{2}\cdots{\bf e}_{2k+2t})({\bf e}_{j_{t+1}-1}\cdots{\bf e}_{j_{p-1}-1})}$
and
$E=\left([2k+2t+1]\setminus\{j_{1},\ldots,j_{t}\}\right)\cup\{j_{t+1}-1,\ldots,j_{p-1}-1\}.$
Then $|E|=2k+p$ and $E\subseteq\mathrm{set}(\beta)$. Furthermore, $\frac{\alpha^{k+p}}{{\bf x}_{F}}={\bf x}_{E}\cdot\beta$, as required.    ∎
@@
(6) $\mathrm{HS}_{i}(I^{s})=\sum_{\begin{subarray}{c}\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}\leq k\leq\left\lfloor\frac{i}{2}\right\rfloor\end{subarray}}\left(\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i-2k\end{subarray}}\frac{\alpha^{i-k}}{{\bf x}_{F}}\cdot I^{s-i+k}\right)$
@@
let p=i2kp=i-2k. Then $k+p\leq\left\lceil\frac{n}{2}\right\rceil-1$ and k+p=ikk+p=i-k. By Lemma 4.13, for any $F\subseteq[n]$ with $|F|=p$, we have $\frac{\alpha^{i-k}}{{\bf x}_{F}}\in\mathrm{HS}_{i}(I^{i-k})$. It follows from [18, Proposition 1.3] that
$\frac{\alpha^{i-k}}{{\bf x}_{F}}\cdot I^{s-i+k}\subseteq\mathrm{HS}_{i}(I^{s}).$
Thus, all terms in the right-hand side are contained in $\mathrm{HS}_{i}(I^{s})$, so $\supseteq$ holds. The converse inclusion follows from Proposition 4.11. This completes the proof.    ∎
As an immediate consequence of Theorem 6, we obtain the generating degree of $\operatorname{HS}_{i}(\mathcal{R}(I))$.
@@
Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal. For all integers $1\leq i<\frac{n}{2}$, the $i$-th homological shift algebra $\operatorname{HS}_{i}(\mathcal{R}(I))$ is generated in degree $\leq i-q$. Here, $q=\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}$.
Let $L_{j}$ denote the path with vertex set $[n]$ and edge set $E(G)\setminus\{j-1,j\}$ for $j=1,\ldots,n$. As before, 0 refers to the vertex $n$. Note that $L_{1},\ldots,L_{n}$ are all the spanning trees of $G$.
Let $I_{j}$ denote the complementary edge ideal of $L_{j}$ for all $j$. By Lemma 1.3, we have $\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})\subseteq\mathrm{HS}_{i}(I^{s})$. It is therefore natural to ask which part of the right-hand of Formula (5) corresponds to $\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})$.
@@
If $i<\frac{n}{2}$, then
$\mathrm{HS}_{i}(I^{s})=\sum_{0\leq k\leq\left\lfloor\frac{i}{2}\right\rfloor}\left(\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i-2k\end{subarray}}\frac{\alpha^{i-k}}{{\bf x}_{F}}\cdot I^{s-i+k}\right).$
We claim that $\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})$ corresponds to the $k=0$ component of the expression above. More precisely,
$\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})=\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i\end{subarray}}\frac{\alpha^{i}}{{\bf x}_{F}}\cdot I^{s-i}\quad\text{for all }s\geq i.$
In fact, by Theorem 4.3,
$\mathrm{HS}_{i}(I_{1}^{s})=\sum_{\begin{subarray}{c}F\subseteq[n]\setminus\{1,n\}\\ |F|=i\end{subarray}}\frac{\alpha^{i}}{\mathbf{x}_{F}}I^{s-i},$
@@
$\mathrm{HS}_{i}(I_{j}^{s})=\sum_{\begin{subarray}{c}F\subseteq[n]\setminus\{j-1,j\}\\ |F|=i\end{subarray}}\frac{\alpha^{i}}{\mathbf{x}_{F}}I^{s-i}.$
This establishes the inclusion $\subseteq$ in the desired equality. For the converse, take any $F\subseteq[n]$ with $|F|=i$. Since every vertex cover of $G$ contains at least $m$ vertices (and $i<\frac{n}{2}=m$ for even $n$, or $i<\frac{n-1}{2}<m$ for odd $n$), $F$ cannot be a vertex cover of $G$. Thus, there exists some $j_{0}\in[n]$ such that $F\cap\{j_{0}-1,j_{0}\}=\emptyset$, which implies $\frac{\alpha^{i}}{{\bf x}_{F}}\cdot I^{s-i}\subseteq\mathrm{HS}_{i}(I_{j_{0}}^{s})$. This proves the claim.
@@
Proposition 4.17.
Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal.
(1) If $n=2m+1$, then $\mathrm{HS}_{2m}(I^{s})=\alpha^{m}\cdot I^{s-m}$ for $s\geq m$. In particular, $\mathrm{HS}_{2m}(\mathcal{R}(I))$ is generated in degree $m$.
(2) If $n=2m$, then $\mathrm{HS}_{2m-2}(I^{s})=\alpha^{m-1}\cdot I^{s-m+1}$ for $s\geq m-1$. In particular, $\mathrm{HS}_{2m-2}(\mathcal{R}(I))$ is generated in degree m1m-1.
@@
(1) In Formula (5), the range of $k$ is given by

### v5

@@ 5.3. Cycles
We only consider the case when $n=2m+1$ (an odd number), as the proof for the other case is analogous and even simpler. Let $j\in\mathrm{set}(\beta)$. Then $j\in[n-1]$ and $j\in u_{k}$ for some $1\leq k\leq s$. If j=n1=2mj=n-1=2m, then by Lemma 5.7.(3), $\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2m-3},{\bf e}_{2m-1}\}\subseteq\{u_{1},\ldots,u_{s}\}$, which implies $j\in N_{1}(\beta)$. Thus, we may assume $j\leq 2m-1$. If $j=\min(u_{k})$ for some $k\in[s]$, then $j\in\{\min(u_{j}):\ 1\leq j\leq s\}\setminus\{n-1\}$, and we are done.
Suppose now that $j\leq 2m-1$ and $j\neq\min(u_{k})$ for any $k=1,\ldots,s$. By Lemma 5.7, we conclude that if $j=2\ell$ for some $\ell$, then $\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-1}\}\subseteq\{u_{1},\ldots,u_{s}\}$; and if $j=2\ell+1$ for some $\ell$, then $\{{\bf e}_{n},{\bf e}_{2},\ldots,{\bf e}_{2\ell}\}\subseteq\{u_{1},\ldots,u_{s}\}$. Furthermore, since $j\leq 2m-1$, we have: If $j=2\ell+1$, then $\ell\leq m-1$; If $j=2\ell$, then $\ell\leq m-1$. This implies $j\in N_{1}(\beta)\cup N_{2}(\beta)$, so the inclusion $\subseteq$ holds. The converse inclusion follows directly from Lemma 5.7.
@@
By this corollary, it is straightforward to verify that $\mathrm{HS}_{i}(\mathcal{R(}I))$ is generated in degree $\leq n$ for all $i\geq 1$. This result is sharpened in Corollary 5.14.
Hereafter, we focus on computing the homological shift ideals of powers of the complementary edge ideal associated with a cycle graph. Note that if $s<\frac{i}{2}$, then $\mathrm{HS}_{i}(I^{s})=0$, so we only consider the case when $s\geq\frac{i}{2}$. In the rest of this section, we always use the convention that $I^{t}=0$ for all $t<0$.
@@
Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal. For $1\leq i\leq n-1$ and $s\geq\frac{i}{2}$, the following inclusion holds:
@@
Write $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$ with $u_{j}\in\mathcal{G}(I(G))$ for all $j=1,\ldots,s$. Set $T:=\{u_{1},\ldots,u_{s}\}$. Denote
@@
By convention, $\max\emptyset=$-$\infty$. By Lemma 5.8, $N_{1}(\beta)\cup N_{2}(\beta)=[2\ell_{1}]\cup[2\ell_{2}+1]$. We consider two cases:
@@
(2) $Q_{1}:=Q\cap\left([2\ell_{1}]\cup[2\ell_{2}+1]\right)\neq\emptyset$: In this case, we can decompose $Q$ as
@@
where the edges $v_{1},\ldots,v_{q},f_{1},\ldots,f_{r}$ are pairwise disjoint, and $j_{t}\in f_{t}$ for $t=1,\ldots,r$. Note that $2q+r=p$ and $q+r\leq\ell_{2}$. For each $t=1,\ldots,r$, let $i_{t}\in[2\ell_{2}+1]$ be such that $f_{t}=x_{j_{t}}x_{i_{t}}$.
Since $\{v_{1},\ldots,v_{q},f_{1},\ldots,f_{r},{\bf e}_{j_{p+1}},\ldots,{\bf e}_{j_{i}}\}\subseteq\{u_{1},\ldots,u_{s}\}$, we may assume, up to reindexing the elements $u_{1},\dots,u_{s}$, that
@@
Let us set $F:=\{i_{1},\ldots,i_{r},j_{p+1}+1,\ldots,j_{i}+1\}$. Then $F$ is an (iq)(i-q)-subset of $[n]$. With this notation, we have
${\bf x}_{Q}\beta=\frac{\alpha^{i-q}}{(x_{i_{1}}\cdots x_{i_{r}})\cdot(x_{j_{p+1}+1}\cdots x_{j_{i}+1})}\cdot\frac{\alpha^{s-i+q}}{u_{1}\cdots u_{s-i+q}}\in\frac{\alpha^{i-q}}{{\bf x}_{F}}I^{s-i+q},$
which implies that ${\bf x}_{Q}\beta$ lies in the component corresponding to $k=q$ on the right-hand side of (4).
The proof for $\ell_{1}>\ell_{2}$ is analogous, so the argument is complete.
In general, the converse of the inclusion (4) fails to hold. For the converse inclusion to hold, it is necessary to remove certain components from the right-hand side of (4). This is accomplished by proving that any subset $F\subseteq\operatorname{set}{\left(\frac{\alpha^{s}}{u_{1}\ldots u_{s}}\right)}$ with $|F|=\left\lceil\frac{n}{2}\right\rceil+k-1$ contains at least $k$ pairwise disjoint edges among $u_{1},\ldots,u_{s}$.
Note that for any integer $k$, $k\leq\left\lfloor\frac{i}{2}\right\rfloor$ if and only if $k\leq\frac{i}{2}$, and $k\geq\left\lceil\frac{i}{2}\right\rceil$ if and only if $k\geq\frac{i}{2}$.
@@
Set $q:=\max\left\{i-\lceil\frac{n}{2}\rceil+1,0\right\}$. If $i\leq\lceil\frac{n}{2}\rceil-1$, then $q=0$, so the right-hand side of (5) coincides with the left-hand side of (4), and the result follows immediately from Lemma 5.10. For $i>\lceil\frac{n}{2}\rceil-1$, we focus on the case $n=2m$ since the argument for $n=2m+1$ is analogous. Since q=im+1q=i-m+1 in this case, we have i=m1+qi=m-1+q. Take $\beta\in\mathcal{G}(I^{s})$ with $i\leq|\mathrm{set}(\beta)|$, and let $Q\subseteq\mathrm{set}(\beta)$ be a subset of size $i$; it suffices to show that $\mathbf{x}_{Q}\beta$ lies in the right-hand side of (5).
To begin with, we write $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$, where $u_{j}\in\mathcal{G}(I(G))$ for all $j=1,\ldots,s$, and set $T:=\{u_{1},\ldots,u_{s}\}$.
@@
$\mathrm{set}(\beta)=[2\ell_{1}]\cup\left(\{\min(u_{j}):\ j=1,\ldots,s\}\setminus\{n-1\}\right).$
Note that the edges ${\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2m-3}$ are pairwise disjoint, and their union is the set [2m2][2m-2]. Since $Q\subseteq\mathrm{set}(\beta)\subseteq[2m-2]$ and |Q|=m1+q|Q|=m-1+q, by the pigeonhole principle, there are $k\geq q$ edges among ${\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2m-3}$ that are entirely contained in $Q$. Let these $k$ edges be ${\bf e}_{2j_{1}-1},{\bf e}_{2j_{2}-1},\ldots,{\bf e}_{2j_{k}-1}$ with $j_{1}<j_{2}<\cdots<j_{k}$.
@@
$\mathrm{set}(\beta)=[2\ell_{2}+1]\cup\left(\{\min(u_{j}):\ j=1,\ldots,s\}\setminus\{n-1\}\right).$
@@
Hence, ${\bf x}_{Q}\beta$ belongs to $k=q$ component of the right-hand side of (5). This completes the proof.
To obtain the converse inclusion, we need the following lemma.
@@
Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal. Suppose $n=2m$ or $n=2m+1$. Given a subset $F\subseteq[n]$ with $|F|=p$ and an integer $k\geq 0$ satisfying $k+p<\frac{n}{2}$, set $i:=2k+p$. Then there exist pairwise distinct edges $u_{1},\ldots,u_{k+p}$ and a subset $E\subseteq\mathrm{set}\left(\frac{\alpha^{k+p}}{u_{1}\cdots u_{k+p}}\right)$ with $|E|=i$ such that
$\frac{\alpha^{k+p}}{\mathbf{x}_{F}}=\mathbf{x}_{E}\cdot\frac{\alpha^{k+p}}{u_{1}\cdots u_{k+p}}\in\mathrm{HS}_{i}(I^{k+p}).$
@@
We only consider the case $n=2m$; the case $n=2m+1$ is analogous. Write $F=\{j_{1}<j_{2}<\cdots<j_{p}\}$. We analyze two cases:
(1) $n\notin F$: Define
$t:=\max\left\{a:j_{1}\leq 2k+2,\ j_{2}\leq 2k+4,\ \ldots,\ j_{a}\leq 2k+2a\right\}.$
By convention, set $t:=0$ when $j_{1}>2k+2$. Then $1\leq j_{1}<\cdots<j_{t}\leq 2k+2t\leq 2m-2$ and $j_{p}>\cdots>j_{t+1}>2k+2t+2$. Set
$\beta:=\frac{\alpha^{k+p}}{(\mathbf{e}_{1}\mathbf{e}_{3}\cdots\mathbf{e}_{2k+2t-1})(\mathbf{e}_{j_{t+1}-1}\cdots\mathbf{e}_{j_{p}-1})},$
and $E:=([2k+2t]\setminus\{j_{1},\ldots,j_{t}\})\sqcup\{j_{t+1}-1,\ldots,j_{p}-1\}.$ By Lemma 5.8, $E\subseteq\mathrm{set}(\beta)$. Moreover, $|E|=2k+p=i$, and $\mathbf{x}_{E}\cdot\beta=\frac{\alpha^{k+p}}{\mathbf{x}_{F}}\in\mathrm{HS}_{i}(I^{k+p})$, as required.
(2) $n\in F$: Define $t:=\max\left\{a:j_{1}\leq 2k+3,\ j_{2}\leq 2k+5,\ \ldots,\ j_{a}\leq 2k+2a+1\right\},$ and set $t:=0$ when $j_{1}>2k+3$. Then $0\leq t\leq p-1$, $\{j_{1},\ldots,j_{t}\}\subseteq[2k+2t+1]\subseteq[2m-1]$, and $j_{t+1}>2k+2t+3$. Set
$\beta:=\frac{\alpha^{k+p}}{(\mathbf{e}_{0}\mathbf{e}_{2}\cdots\mathbf{e}_{2k+2t})(\mathbf{e}_{j_{t+1}-1}\cdots\mathbf{e}_{j_{p-1}-1})}.$
and $E:=\left([2k+2t+1]\setminus\{j_{1},\ldots,j_{t}\}\right)\sqcup\{j_{t+1}-1,\ldots,j_{p-1}-1\}.$ Then $|E|=2k+p=i$, and by Lemma 5.8, $E\subseteq\mathrm{set}(\beta)$. It follows that $\frac{\alpha^{k+p}}{\mathbf{x}_{F}}=\mathbf{x}_{E}\cdot\beta\in\mathrm{HS}_{i}(I^{k+p})$, as required.
@@
(6) $\mathrm{HS}_{i}(I^{s})=\sum_{\begin{subarray}{c}\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}\leq k\leq\left\lfloor\frac{i}{2}\right\rfloor\end{subarray}}\left(\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i-2k\end{subarray}}\frac{\alpha^{i-k}}{\mathbf{x}_{F}}\cdot I^{s-i+k}\right).$
@@
let p=i2kp=i-2k. Then k+p=ikk+p=i-k, and it follows that $k+p\leq\left\lceil\frac{n}{2}\right\rceil-1$. By Lemma 5.12, for any $F\subseteq[n]$ with $|F|=p$, we have $\frac{\alpha^{i-k}}{\mathbf{x}_{F}}\in\mathrm{HS}_{i}(I^{i-k})$. Applying [15, Proposition 1.3], we deduce that
$\frac{\alpha^{i-k}}{\mathbf{x}_{F}}\cdot I^{s-i+k}\subseteq\mathrm{HS}_{i}(I^{s}).$
Thus, every term in the right-hand side is contained in $\mathrm{HS}_{i}(I^{s})$, whence the inclusion $\supseteq$ holds. The reverse inclusion follows immediately from Proposition 5.10. This completes the proof.
We remark that Equality (6) still holds even when $s<\left\lfloor\frac{i}{2}\right\rfloor$. In this case, $\mathrm{HS}_{i}(I^{s})=0$ by Theorem 4.4, and the right-hand side of this equality is also zero, since si+k<0s-i+k<0 for all $k<\left\lfloor\frac{i}{2}\right\rfloor$.
Theorem 5.13 yields the generating degree of $\mathrm{HS}_{i}(\mathcal{R}(I))$.
@@
Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal. For all integers $1\leq i<\frac{n}{2}$, the $i$-th homological shift algebra $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated in degree $\leq i-q$, where $q=\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}$.
Proof.
It is straightforward to verify that $\mathrm{HS}_{i}(I^{s})=I^{s-i+q}\cdot\mathrm{HS}_{i}(I^{i-q})$ for any $s\geq i-q$. The desired conclusion then follows directly from this identity.
@@
Let $G$ be a cycle of length $n$. For each $j=1,\ldots,n$, let $L_{j}$ denote the path graph with vertex set $[n]$, obtained by deleting the edge $\{j-1,j\}$ from $G$. Here, we adopt the convention that vertex 0 corresponds to vertex $n$ (so the edge deleted for $L_{1}$ is $\{n,1\}$). Note that $L_{1},\ldots,L_{n}$ are exactly all the spanning trees of $G$. For each $j$, let $I_{j}$ denote the complementary edge ideal of $L_{j}$. By Lemma 2.5, we immediately obtain the inclusion
$\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})\subseteq\mathrm{HS}_{i}(I^{s}).$
A natural question then arises: Which component of the decomposition of $\mathrm{HS}_{i}(I^{s})$ given in Formula (6) does the sum $\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})$ correspond to?
Suppose $1\leq i<\frac{n}{2}$. For this range of $i$, $\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}=0$, since $i<\frac{n}{2}\leq\left\lceil\frac{n}{2}\right\rceil$ implies $i-\left\lceil\frac{n}{2}\right\rceil+1<1$. Theorem 5.13 thus simplifies to
$\mathrm{HS}_{i}(I^{s})=\sum_{0\leq k\leq\left\lfloor\frac{i}{2}\right\rfloor}\left(\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i-2k\end{subarray}}\frac{\alpha^{i-k}}{\mathbf{x}_{F}}\cdot I^{s-i+k}\right).$
We claim that $\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})$ is contained in the $k=0$ component of the above decomposition. In other words, the following inclusion holds for all $s\geq i$:
$\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})\subseteq\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i\end{subarray}}\frac{\alpha^{i}}{\mathbf{x}_{F}}\cdot I^{s-i}.$
Moreover, this inclusion becomes an equality when $s=i$.
First, we verify the inclusion $\subseteq$. By Theorem 5.2, $\mathrm{HS}_{i}(I_{j}^{s})$ has the following decomposition:
$\mathrm{HS}_{i}(I_{1}^{s})=\sum_{\begin{subarray}{c}F\subseteq[n]\setminus\{1,n\}\\ |F|=i\end{subarray}}\frac{\alpha^{i}}{\mathbf{x}_{F}}I_{1}^{s-i},$
@@
$\mathrm{HS}_{i}(I_{j}^{s})=\sum_{\begin{subarray}{c}F\subseteq[n]\setminus\{j-1,j\}\\ |F|=i\end{subarray}}\frac{\alpha^{i}}{\mathbf{x}_{F}}I_{j}^{s-i}.$
Since $I_{j}\subseteq I$ for each $j$, we have $I_{j}^{s-i}\subseteq I^{s-i}$. This implies that every term in $\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})$ is contained in the $k=0$ component of $\mathrm{HS}_{i}(I^{s})$, proving the inclusion $\subseteq$.
Next, we establish the reverse inclusion $\supseteq$ in the case $s=i$. Recall that a vertex cover of a graph $G$ is a subset $C\subseteq V(G)$ such that every edge of $G$ is incident to at least one vertex in $C$. Since $i<\frac{n}{2}$ and the minimum vertex cover of a cycle $G$ has size $\left\lceil\frac{n}{2}\right\rceil$, $F$ cannot be a vertex cover of $G$. By definition of vertex covers, there exists at least one edge $\{j_{0}-1,j_{0}\}$ of $G$ such that $F\cap\{j_{0}-1,j_{0}\}=\emptyset$. It follows that $F\subseteq[n]\setminus\{j_{0}-1,j_{0}\}$, so
$\left(\frac{\alpha^{i}}{\mathbf{x}_{F}}\right)\subset\mathrm{HS}_{i}(I_{j_{0}}^{i}).$
Thus, every term in the $k=0$ component (when $s=i$) is contained in $\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{i})$, proving the inclusion $\supseteq$. Combining both inclusions, the claim is established.
@@
Corollary 5.16.
Let $G$ be a cycle of length $n$ and $I$ denote its complementary edge ideal.
(1) If $n=2m+1$, then $\mathrm{HS}_{2m}(I^{s})=\alpha^{m}\cdot I^{s-m}$ for all integers $s\geq m$. In particular, $\mathrm{HS}_{2m}(\mathcal{R}(I))$ is generated in degree $m$.
(2) If $n=2m$, then $\mathrm{HS}_{2m-2}(I^{s})=\alpha^{m-1}\cdot I^{s-m+1}$ for all integers $s\geq m-1$. In particular, $\mathrm{HS}_{2m-2}(\mathcal{R}(I))$ is generated in degree m1m-1.
@@
(1) In Formula (6) of Theorem 5.13, the range of $k$ is given by